OpenStudy (anonymous):

The parametric equations for a line L1 are as follows: x = −5−6t y = 5+9t z = −2+3t Let L2 be the line parallel to L1 and passing through the point (4, −3, −3). Find the point P on L2 whose x-coordinate is −12.

OpenStudy (tkhunny):

Have you considered: x = 3−6t y = -3+9t z = −3+3t ?? Now what?